OpenStudy (sammy_22):

What is the angle of 135 degrees equal to in radian measure?

OpenStudy (mww):

Make sure you know the conversion formula For me I always remember one thing \[l = r \theta\] \[When ~ l=C = 2\pi r = r \theta, ~ \theta = 2\pi = 360^{0} ~so~ \pi ~rad = 180^0\] We always start with pi radians = 180 degrees. To change from 180 to 135 degrees, divide both sides by 180 and times by 135 thus 135 degrees = pi/180 x 135 In general \[x^0 = \frac{ x \pi }{ 180 } rad\]

OpenStudy (mathmale):

True, \[\pi ~rad=180~\deg\]

OpenStudy (mathmale):

From this we could derive the proper conversion factor:\[\frac{ \pi ~rad }{ 180~ \deg}\]

OpenStudy (mathmale):

Now multiply 135 deg by this conversion factor. Note how "deg" will cancel out, leaving you with a result expressed in radians, as desired.